How many liters of a 10% silver iodide solution must be mixed with 7L of a 4% silver iodide solution to get a 6% solution?
Answer by stanbon(57246)

How many liters of a 10% silver iodide solution must be mixed with 7L of a 4% silver iodide solution to get a 6% solution?
---------------------
Equation:
silver + silver = silver
0.10x + 0.04*7 = 0.06(x+7)
----
Multiply thru by 100 to get:
10x + 4*7 = 6x + 6*7
4x = 2*7
x = 3.5L (amt. of 10% needed)
=================================
Cheers,
Stan H.
==========